Astronomy Report – February 11th

Astronomy Report

Name: Ghanim Alotaibi

Crew: 221
Date: 11 Feb 2020

One image was obtained last night for V0960 Mon. Crew astronomer tried to perform a photometry measurement but was unable to do so.


Robotic Telescope Requested MDRS-14

Objects to be Imaged this Evening:

  1. V CAS (Resubmitted last night)
  2. V0960 Mon (Resubmitted today)

Images submitted with this report: seeing profile of V0960 MON

Problems Encountered: As seen from the seeing profile, AstroImage j was unable to differentiate the target star from the background noise, however, it was possible to differentiate the star visually. This might be because the target star is too faint. The focusing of the image overall looks good and the seeing profile for random stars also looks good.

GreenHab Report – February 11th

Crew 221 GreenHab Report 11FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on. Door opened today from 12:15 pm to 4:06 pm.

Average temperatures: 22°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 15.85 gallons

Daily water usage for research and/or other purposes: 0.53 gallon

Water in Blue Tank: 178.8 gallons

Time(s) of watering for crops: 2

Changes to crops: We have a cucumber! The zucchinis and cucumbers are flowering, and the strawberry plant has begun growing a strawberry. The tips of the narcissus are incredibly dry and browning. They have been watered regularly so I’m concerned the heater is drying them out. Should we move them to a lower level?

Narrative: Today was exciting as we harvested quite a bit from the GreenHab on Shannon’s recommendation for tonight’s dinner. Rosemary potatoes and salad! We are feasting like kings! The plants were watered this morning and this evening, and excitingly one of our Blue corn plants in Martian soil has begun to sprout. That’s right; Crew 221 has officially grown crops on Mars.

Harvest: 71g Cucumber, 20g lettuce, 6g butter lettuce, 14g onions, 4g rosemary.

Support/supplies needed: NA

Other Notes: Myself and the Crew Engineer noticed the water level in the GreenHab tank is much lower than expected. The water has just dropped below the level of the upper tap, meaning it is around 1/3 full (~100 gallons). We have been basing our water amount each day on subtracting the amount used every day from a full tank at the start of Sol 0. Due to the busy schedule of Crew 220, and the repairs being done by David and Atila, Crew 221 did not complete all the water refilling of the station prior to the start of our mission, and the GreenHab tank had to be refilled by Atila on Sol 1. We were told the GreenHab tank was filled to full (300 gallons), however we have kept a careful record of our water usage, and while our numbers tell us we should have 178.8 gallons, this is clearly not the case. We should have sufficient water in the GreenHab to get us to the end of our mission, however it will be cutting it close.

Science Report – February 11th

Crew 221 Science Report 11-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we completed our sample catalogue: we identified the probable bacteria species for each sample that showed growth. One of the interesting samples was the bacteria growth from a sample taken from the Candor Chasma canyon floor. The bacteria is called Flavobacterium Capsolatum.

2. ISRU project
Today, corn sprouted from the half earth soil half grey martian soil today!

3. Mung Beans

Nothing Further to report.

4. Radio Communication

Today, we disassembled the radio and Crew Engineer Louis and Crew Astronomer Ghanim reassembled the radio on an EVA, proving that it is compatibly as a mobile communication device!

Journalist Report – February 11th

Crew 221 Journalist Report 11Feb2020

Author: Hemani Kalucha, Crew Journalist

Sol 9

Title: Corn on Mars (with a little bit of help)

This morning, we tested radio operations outside the hab. Crew Engineer Louis had already successfully built an engineer and got it working inside the hab over the last few days. The test then involved seeing if it could be reassembled with ease on an EVA in a short period of time. It was exciting to see that the radio could be reassembled in under an hour on EVA in two bulky astronaut suits. The second EVA was supposed to test if a team could reassembled the radio with no prior knowledge of radios (this team was Commander Grulich, Hemani, and Jess), however, due to a typo error in the request, the EVA had to be cancelled. We hope to recover this EVA tomorrow. Meanwhile, the rover was successfully built, but the team ran into several issues while trying to program it, and will try again to get the rover running tomorrow. On a happier note, the team spent the day filming various activities, which they will use in outreach videos, and also took some awesome astronaut profile pictures. Also for dinner tonight, we have the luxury of eating tons of leafy greens, all harvested from our very own greenhouse! The sunny, warm days on Mars are keeping our spirits up as we edge closer to the end of the mission!

EVA Report – February 11th

Crew 221 EVA # 10 Report 11 Feb 2020

EVA # 10

Author: Luis Monge (Engineer)

Purpose of EVA: We will be performing the deployment of a dipole antenna and document the whole process. The antenna and required implements will be carried in the rovers and assembled close to them at the destination.

Start time: 10:00 am

End time: 11:30 pm

Narrative: Today we conducted the first of the two EVA for antenna deployment exercises. L.Monge, and G.Alloitabi went out to drive towards Cowboy Corner with the aim to practice long wave antenna installation. The Martian weather was sunny and chilly during the EVA. L.Monge took the lead and drove directly to the spot, once there proceeded to assemble the antenna.

The procedure was documented with video and pictures, the antenna was disassembled the components stored in the rovers and the crew headed back home on time.

Destination: Cowboy Corner

Coordinates (use UTM NAD27 CONUS): 4252600/518770 (Cowboy Corner)

EVA Participants: Luis, Ghanim

Road(s) and routes per MDRS Map: Road(s) and routes per MDRS Map: Park along Cow Dung Road (0110) to stop at Cowboy Corner

Mode of travel: Driving and Walking

Vehicles you will be using (If applicable): Curiosity and Spirit

Operations Report – February 11th

Crew 221 Operations Report 11Feb2020

SOL: 9

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 7:40 pm)

Hours run: 13:00

From what time last night: 6:00pm

To what time this morning: 7:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 78%

Diesel Reading – 85%

Station Propane Reading – 54%

Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 215 gals

Static to Loft Pump used – yes

Water in GreenHab – 178.8 gals (water is probably lower than estimated, please refer to greenhab report)

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 130.4

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 51%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 136.6

Beginning charge: 100%

Ending charge: 31%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues, now watered with 2l/day. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 1 incubator runing, microbiology experiments going on and cultures are bieng grown in agar plates.

Summary of RAM operations: robot assembly almost completed and expected to continue, radio comunication experiment continued today.

Summary of any observatory issues: nominal

Summary of health and safety issues: crew memeber with flu continues recovery.

Questions, concerns and requests to Mission Support: a mouse trap is installed at the kitchen, as we suspect additional mouse is at the kitchen.

Sol Summary – February 11th


Summary Title: Wish List and Microbiology Sample Catalogue

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol summary: The crew had a lovely Canadian/Indian Cultural night, learning about two cultures and getting our names written in Hindi.

Today we woke up and wrote our wish list for the activities we would love to do with our remaining time on Mars. Number one is movie night, games night and salsa night as we keep on postponing it due to the high number of projects we would like to achieve. We still have 4 days and only one cultural night left.

The first EVA in the morning left just 15 min late as one of the wishes from the commander was to take the EVA checklists more seriously. Still the EVA Support team forgot to put a camera in the airlock, which caused the EVA team to do the depressurization twice.

Once on the road, the EVA team managed to install the antenna on the desired location and achieved all their objectives. The rest of the crew was working on finishing the robot and managed to see it turning on for the first time. Unfortunately, the crew could not manage to connect to the robot and has to troubleshoot further tomorrow. For lunch, the team had tortilla bowls with fresh mung bean sprouts that the crew grew over the last couple of days. Mung beans are a powerhouse of vitamins such as A,B,C and E as well as minerals calcium, magnesium, potassium and phosphorous. It also contains zinc, chromium, iron and they taste very fresh and like peas, which is great for Martians in confined space.

Also, today the inherited Kambucha batch of crew 220 should be ready to be drunk by the crew. A first try will be done tomorrow.

During the afternoon, the EVA team got delayed one hour as the crew only had four EVA radios and two of them had been used before and were not recharged in time. In addition, the crew made a mistake in the EVA window and so when the team was finally ready to go, the team was informed that the EVA window is already over. Therefore, the EVA was canceled.

The free time was used to take four crew members’ profile pictures, work on the sample catalogue of the microbiology samples and identify the bacteria based on their appearance such as, form, elevation, surface, opacity and color. The samples from February 3rd were then discarded to avoid any bad bacteria growth. As we are working on fulfilling our wish list, we will have a games night tonight, after a lovely dinner with fresh salad and cucumber from the Greenhab harvested by J. Todd!

Ad Astra!

Anomalies in work: None

Update mouse in hab: Mouse trab reinstalled no new mouse caught

Weather: sunny no winds

Crew Physical Status: Crew status with signs of flu feel better, also crew with knee issues is feeling better.

Reports to be filed:

Sol Summary

Photo Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

EVA Request

EVA Report 10

Journalist Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site